OpenStudy (anonymous):

Integrate: (x^2 e^3x+x+3/ x^2)

OpenStudy (anonymous):

\[\int\left(x^2e^{3x}+x+\dfrac{3}{x^2}\right)~dx\] or \[\int\frac{x^2e^{3x}+x+3}{x^2}~dx~~?\]